OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (-13,-14) and (-20,-15)

OpenStudy (anonymous):

The slope formula (given two points \((x_1,y_1)\) and \((x_2,y_2)\) is: \[Slope = \frac{y_2-y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

So to find it in this example you have: \[Slope = \frac{-15- (-14)}{-20 - (-13)}\]

OpenStudy (anonymous):

-1/-7?

OpenStudy (anonymous):

1/7

OpenStudy (anonymous):

-1/-7 = 1/7 yes.

OpenStudy (anonymous):

something is wrong with server, I'm not deleting anything pol

OpenStudy (anonymous):

I didn't say you were.

OpenStudy (anonymous):

well, it's so tedious job